OpenStudy (anonymous):

every month, tristen deposits $488 into an interest-bearing account to save for a down payment on a house. The interest rate on the account is 5.27% compounding annually. What is the present value of the investment if tristan purchases a house in 15 years? a)$4,974.08 b)$5,236.22 c)$59,689.00 d)$62,834.61 @Hunus

OpenStudy (hunus):

\[FV = \frac{d*((1+\frac{r}{n})^{nt} - 1)}{\frac{r}{n}}\]

OpenStudy (hunus):

That's the formula you will need to use. d represents the amount she deposits each time period

OpenStudy (anonymous):

thanks

OpenStudy (hunus):

yup :)

OpenStudy (anonymous):

i did that formula but i keep getting 21,891.47...its not one of the choices

OpenStudy (anonymous):

@e.mccormick

OpenStudy (e.mccormick):

OK, so you need the future value after 15 years of that monthy.

OpenStudy (anonymous):

yeah but how would i find it?

OpenStudy (e.mccormick):

I wanted to check that formula, I think it is right, but just to be sure.... looking keeps getting me the wrong ones. LOL. I had it not long ago.

OpenStudy (anonymous):

ok lol

OpenStudy (e.mccormick):

Ah, I see a few that divide by the interest, but when you divide by a fraction, it inverts and multiplies, so that makes sense.

OpenStudy (e.mccormick):

So what numbers did you put in hwere?

OpenStudy (e.mccormick):

AH HA! I see an issue with how this formula works.... This formula, as I recall, it based on the payment being the same as the compounding period. In this problem, the compounding period is annually, but the depostis are monthy. You need to do 12 months of deposits as a single annual one.

OpenStudy (anonymous):

d- 488 r/n - 5.27%/2=0.02635 nt- 15*2=30

OpenStudy (anonymous):

so instead of 2 do i put 12?

OpenStudy (e.mccormick):

You may be able to get it to work that way. It would break up the interest, but I am not sure if that will work right. I am still looking for the formula I used last week. LOL. If I find it, I'll run it with a payment of 5856 as if it was one payment a year.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

i tried it n it didnt work...

OpenStudy (e.mccormick):

\[A = P * \left[\frac{(1+i)^n - 1 }{i}\right] * (1+i)\] (if payment are being made at the start) (it will be P * [{(1+i)^n - 1 }/i] if payments are made at the end of the year) Where : A = final amount P = installment each time n = total number of installments i = interest rate for that tenure (example if yearly return is 24% , but payments are made monthly then i = 24/12 = 2%) That seems to be the same thing as above....

OpenStudy (anonymous):

i got 982.44?

OpenStudy (e.mccormick):

There is a problem with the question. Lets assume matress method. Every month, 488 under a matress. So 488*12*15=? So no interest at all.

OpenStudy (e.mccormick):

Look at that answer next to your possible answers and you will see the issue.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its too high

OpenStudy (e.mccormick):

And that is without interest... They messed up the problem.

OpenStudy (anonymous):

but if you multiply that by 5.27% it is closer...

OpenStudy (e.mccormick):

No, it does not work that way.

OpenStudy (anonymous):

oh ok

OpenStudy (e.mccormick):

The answer should be well over $100,000 with that rate and deposit system. Just the depostired money is close to $90,000 so the investment would have compounded to far more than that.

OpenStudy (e.mccormick):

Check the numbers in the problem. make sure it is 15 years and $488.

OpenStudy (anonymous):

it is

OpenStudy (e.mccormick):

And it says present value, not return, or interest value?

OpenStudy (e.mccormick):

hmmm.... even return on investment would not get one of those numbers.

OpenStudy (anonymous):

its present value

OpenStudy (e.mccormick):

You have the unfortunate situation of being the recipient of a set of answers that are male bovine animal's experiment. They make no sense at all.

OpenStudy (e.mccormick):

oops... excrement... You get the point.

OpenStudy (anonymous):

ya

OpenStudy (e.mccormick):

The other problem I see is they even used the wrong term. "present value of the investment" is a reference to how things decrease in buying power due to inflation. So, i don't know why they would word the poroblem as a future value problem then say it was a present value problem.

OpenStudy (anonymous):

The problem isn't wrong... They just want you to find out the present value NOT what he's going to have in 15 years. I hate this problem, I picked $4,974.08. I don't know if I'm right or not I just know the question isn't wrong.